OpenStudy (anonymous):

Assume that 1,000 scores are normally distributed with a mean of 72 and a standard deviation of 4. How many scores can we expect to be between 75 and 82?

OpenStudy (anonymous):

z-score for 75 = (75-72)/4 = 3/4 = .75 z-score for 82 = (82-72)/4 = 10/4 = 2.5

OpenStudy (anonymous):

use a table such as http://training.ce.washington.edu/pgi/Modules/08_specifications_qa/Images/main_pictures/normal_table.gif to find the area between them

OpenStudy (anonymous):

actually that table sucks

OpenStudy (anonymous):

P( .75 < Z < 2.5) = P(Z < 2.5) - P(Z < .75) = .9938 - .7734 = .2204

OpenStudy (anonymous):

thansk man